OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 69. n + (n + 2) = 69 n + (n + 1) = 69 n − (n + 1) = 69 n + 2n = 69

OpenStudy (anonymous):

do you know what consecutive integers are?

OpenStudy (anonymous):

no equation for this, just make an educated guess 32 + 33 = 65 too small, try again

OpenStudy (anonymous):

try the last one ;P solve for n first! :D

OpenStudy (anonymous):

There is one and it is one of the choices. buzzz

OpenStudy (anonymous):

hahaha @ChmE =P and heres my hint :) Hint: Use the last choice and solve for n :P

OpenStudy (anonymous):

i got it its n+(n+2)=39 so the two consectuive intergers would be 33 and 34 and this equals 69

OpenStudy (anonymous):

does 33+34 = 69 ???????????????????? no

OpenStudy (anonymous):

not quite... here follow this :)